OpenStudy (anonymous):

rename each fraction as a decimal 3/5 and 6/8 grade five level.

OpenStudy (anonymous):

either divide or write each one with a denominator that is a power of ten, like 10 , 100, 1000, etc

OpenStudy (anonymous):

for the first one you could write \[\frac{3}{5}=\frac{6}{10}=.6\] or you could divide 3 by 5 using long division

OpenStudy (anonymous):

second one probably easiest to reduce the fraction first\[\frac{6}{8}=\frac{3}{4}\] which you probably recognize as .75